Shift report 20030321 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Fri Mar 21 2003 - 15:47:37 EST

  • Next message: Apache: "Shift report 20030321 16:00-24:00"
    ------------------------------------------------------------
    Date: Fri Mar 21 15:47:36 2003, Shift: 08:00-16:00
    Supervisor: Eun-Joo Kim, with: 
    Log book#14 pages 70 - 70
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam was dumped around 10am, and back to 11am.
    
    2. Data Run : FS(12-deg, 1/4A), MRS(40-deg, 2000B)
        - Run 8569, 8570, 8571, 8572, 8573
        - 13k trigger2, 608k trigger3     
    
       Run8560 : D1,D2(1/4A), D3,D4(zero-field)
                 0.7k trigger 2
                 We need to take zero-field run again...
    
       Run 8561 : Camac pedestal
       Run 8565 : TPC Pedestal
       Run 8568 : DVM 
    
       Sum of triggers at current Setting, so far
       Run 8554-8573 :  31k trigger2  
       Run 8557-8573 : 919k trigger3
      
       Plan to do
       - zero field run at the end of store
       - more 19k trigger2 at FS 12degree, 1/4A
      
    3. No new detector problems observed. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Fri Mar 21 2003 - 15:51:47 EST